OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

(x-2)2+(y+1)2=64

OpenStudy (anonymous):

The equation of a circle is (x - h)2 + (y - k)2 = r2 where (h, k) represents the circle's center and r represents the circle's radius.

OpenStudy (anonymous):

Since -1 is negative though you will change the negative to a positive sign

OpenStudy (amistre64):

a circle centered at the origin with a given radius (r) can be equated as: x^2 + y^2 = r^2 when we want to move this about, we introduce soem extra parts to be able to adjust this with. (x-cx)^2 + (y-cy)^2 = r^2 where (cx,cy) is the center of the circle